### v1

@@ 2.1.1 Point Estimation, MSE Expansions, and Bandwidth Choices
  • Given a (preliminary) bandwidth, the package rd2d implements several variance estimators for each group $t\in\{0,1\}$: $V_{t,\mathbf{x}}$ is replaced by $\widehat{V}_{t,\mathbf{x}}=\mathbf{e}_{1}^{\top}\widehat{\boldsymbol{\Gamma}}_{t,\mathbf{x}}^{-1}\widehat{\boldsymbol{\Sigma}}_{t,\mathbf{x},\mathbf{x}}\widehat{\boldsymbol{\Gamma}}_{t,\mathbf{x}}^{-1}\mathbf{e}_{1}$, where $\widehat{\boldsymbol{\Sigma}}_{t,\mathbf{x},\mathbf{x}}$ is either a heteroskedasticity-consistent (HC) or cluster-consistent (CR) variance estimator based on replacing the unknown residuals $\varepsilon_{i,t,\mathbf{x}}$ with the plug-in residuals estimates $\widehat{\varepsilon}_{i,t,\mathbf{x}}=Y_{i}-\mathbf{R}_{p}(\mathbf{X}_{i}-\mathbf{x})^{\top}\widehat{\boldsymbol{\beta}}_{t}(\mathbf{x})$. More precisely, the package allows for the following options:
@@
      Construct an MSE-optimal bandwidth choice for estimating the linear combination $\sum_{|\mathbf{k}|=p+1}\frac{\mu_{t}^{(\mathbf{k})}(\mathbf{x})}{\mathbf{k}!}\mathbf{e}_{1}^{\top}\widehat{\boldsymbol{\Gamma}}_{t,\mathbf{x}}^{-1}\widehat{\boldsymbol{\vartheta}}_{t,\mathbf{x}}(\mathbf{k})$, using a $q$th order local polynomial estimator and the bandwidth $\widehat{c}$ for the coefficients $\mathbf{e}_{1}^{\top}\widehat{\boldsymbol{\Gamma}}_{t,\mathbf{x}}^{-1}\widehat{\boldsymbol{\vartheta}}_{t,\mathbf{x}}(\mathbf{k})$. To implement this bandwidth choice, $\widehat{c}$ is used for variance estimation, and a preliminary nearest-neighbor-based polynomial regression approximation is used for bias estimation. The resulting data-driven bandwidth choice is:
@@ 2.1.2 Statistical Inference
where $\widehat{\tau}(\mathbf{x})$ and $\widehat{V}_{t,\mathbf{x}}$ are constructed using the bandwidth $h$, after choosing the appropriate HC and CR variance estimator, as explained above in the context of bandwidth selection. In practice, the bandwidth is chosen to be (I)MSE-optimal, and thus the sampling distribution of the statistic satisfies the following pointwise distributional approximation:
@@
The standard confidence interval estimator with $100(1-\alpha)\%$ nominal coverage is
@@
where $\Phi_{\alpha}$ is the $\alpha$th quantile of the standard Gaussian distribution. However, for “large” bandwidths such as the (I)MSE-optimal choice, $\operatorname{CI}(\mathbf{x})$ will be invalid due to the bias $\text{Bias}(\mathbf{x})$, thereby delivering empirical coverage well below its nominal target. A solution to this problem is to employ ad-hoc undersmoothing, that is, to implement $\operatorname{CI}(\mathbf{x})$ with a “smaller” bandwidth $h$ relative to the (I)MSE-optimal choice. However, Calonico et al. (2018, 2022) showed that undersmoothing is sub-optimal (possibly invalid) under standard assumptions, while the robust bias-correction (RBC) methodology introduced by Calonico et al. (2014) enjoys validity and better (in some cases optimal) higher-order distributional properties. The core idea behind the RBC approach can be summarized as follows: (i) employ the (I)MSE-optimal bandwidth for constructing the point estimator $\widehat{\tau}(\mathbf{x})$, (ii) de-bias (bias correct) the numerator of the statistic $\operatorname{T}(\mathbf{x})$, and (iii) adjust the variance estimate to account for the variability introduced by the debiasing of $\widehat{\tau}(\mathbf{x})$.
@@
which is constructed with an (I)MSE-optimal bandwidth choice for the $p$th order local polynomial point estimator $\widehat{\tau}(\mathbf{x})$. (This amounts to a form of robust bias correction because $\widehat{\tau}_{q}(\mathbf{x})=\widehat{\tau}(\mathbf{x})-h^{p+1}\widehat{B}_{\mathbf{x}}$, where $\widehat{B}(\mathbf{x})$ is an “estimate” of $B_{\mathbf{x}}$.) As it is customary in practice, the package rd2d employs $p=1$ and $q=2$ as defaults.
@@
where $\mathsf{q}_{\alpha}$ denotes a suitably chosen quantile to control false rejections uniformly over $\mathcal{B}$. In practice, the continuous assignment boundary is discretized to consider the $J$ cutoff in $\mathbf{b}=(\mathbf{b}_{1},\cdots,\mathbf{b}_{J})^{\top}$ jointly. Then, the quantile is defined as
$\mathsf{q}_{\alpha}=\inf\left\{u\geq 0:\ \mathbb{P}\Big{[}\max_{1\leq j\leq J}\big{|}\mathbf{e}_{j}^{\top}\widehat{\mathbf{V}}_{\mathbf{b}}^{1/2}\mathbf{Z}\big{|}\geq u\Big{|}\text{Data}\Big{]}\leq\alpha\right\},$
where the $J$-dimensional standard Gaussian vector $\mathbf{Z}\thicksim\mathsf{N}(\mathbf{0}_{J},\mathbf{I}_{J})$ is independent of the data, and the $(J\times J)$ covariance matrix $\widehat{\mathbf{V}}_{\mathbf{b}}=[\widehat{V}_{\mathbf{b}_{j},\mathbf{b}_{k}}\widehat{V}_{\mathbf{b}_{j},\mathbf{b}_{j}}^{-1/2}\widehat{V}_{\mathbf{b}_{k},\mathbf{b}_{k}}^{-1/2}:1\leq i\leq J,1\leq k\leq J]$ where $\widehat{V}_{\mathbf{b}_{j},\mathbf{b}_{k}}=\widehat{V}_{1,\mathbf{b}_{j},\mathbf{b}_{k}}+\widehat{V}_{0,\mathbf{b}_{j},\mathbf{b}_{k}}$ with
@@
$\boldsymbol{\Sigma}_{t,\mathbf{b}_{j},\mathbf{b}_{k}}=\frac{h^{2}}{n}\sum_{i=1}^{n}\mathbf{r}_{p}\Big{(}\frac{\mathbf{X}_{i}-\mathbf{b}_{j}}{h}\Big{)}\mathbf{r}_{p}\Big{(}\frac{\mathbf{X}_{i}-\mathbf{b}_{k}}{h}\Big{)}^{\top}K_{h}(\mathbf{X}_{i}-\mathbf{b}_{j})K_{h}(\mathbf{X}_{i}-\mathbf{b}_{k})\varepsilon_{i,\mathbf{b}_{j}}\varepsilon_{i,\mathbf{b}_{k}}\mathds{1}(\mathbf{X}_{i}\in\mathcal{A}_{t}),$
for $t\in\{0,1\}$, which is robust to unknown conditional heteroskedasticity. The clustered-robust analogue formula is omitted to save space; see Zeileis (2004) and Zeileis et al. (2020).
@@ 3.1 Function rd2d()
> result.rd2d\$main
> result.rd2d\$main.A0
> result.rd2d\$main.A1
@@
1 0.000 50.000 18.511 18.544 18.511 18.544 775 1650
5 0.000 40.000 16.679 16.709 16.679 16.709 1199 2322
10 0.000 27.500 14.352 14.378 14.352 14.378 1482 2693
15 0.000 15.000 12.690 12.713 12.690 12.713 1403 2381
21 0.000 0.000 16.778 16.808 16.778 16.808 2595 2098
25 10.000 0.000 13.587 13.611 13.587 13.611 1706 2303
30 22.500 0.000 12.206 12.228 12.206 12.228 1236 2095
35 35.000 0.000 12.492 12.514 12.492 12.514 919 1675
40 47.500 0.000 16.748 16.778 16.748 16.778 766 1562
Listing 4: Functionality of rd2d() with output = "bw"
@@ 3.2 Function rdbw2d()
The function rdbw2d() is used for MSE (or IMSE) optimal bandwidth implementation, and is used internally in rd2d() when user does not specify bandwidth choices manually. The function takes the same input data as rd2d(), that is, an outcome vector y, a bivariate location score matrix X, a treatment indicator vector t, and a grid of evaluation points b along the boundary $\mathcal{B}$. In addition, the option bwselect encodes four options of bandwidth type: (i) "mserd" finds the MSE-optimal bandwidth for estimating (derivatives of) treatment effect, (ii) "imserd" finds the integreated MSE optimal bandwidth for estimating (derivatives of) treatment effect, (iii) "msetwo" finds the MSE optimal bandwidth for estimation (the derivatives of) conditional means of two potential outcome variables, (iv) "imsetwo" finds the integrated MSE optimal bandwidth for estimation (the derivatives of) conditional means of two potential outcome variables. The default is bwselect = "mserd". An additional Boolean argument stdvar indicates whether the covariates are first standardized to unit standard deviation in each coordinate, in which case the optimal bandwidth is estimated and then converted back to the original scale. The default is stdvar = FALSE.
@@
1 0.000 50.000 22.719 22.719 22.719 22.719
5 0.000 40.000 21.253 21.253 21.253 21.253
10 0.000 27.500 18.355 18.355 18.355 18.355
15 0.000 15.000 14.514 14.514 14.514 14.514
21 0.000 0.000 21.382 21.382 21.382 21.382
25 10.000 0.000 16.965 16.965 16.965 16.965
30 22.500 0.000 13.772 13.772 13.772 13.772
35 35.000 0.000 14.683 14.683 14.683 14.683
40 47.500 0.000 21.170 21.170 21.170 21.170
Listing 5: Functionality of rdbw2d()
@@
30 22.500 0.000 0.6109 7.3862 0.0000 [0.4338, 0.7469]
35 35.000 0.000 0.5768 6.4730 0.0000 [0.3670, 0.6853]
40 47.500 0.000 0.4654 6.7582 0.0000 [0.3603, 0.6543]
Listing 6: Functionality of rd2d.dist()
@@
Listing 7: Functionality of rd2d.dist() with CBuniform = TRUE
@@
Listing 8: Functionality of rd2d.dist() with output = "bw"
@@
> result.dist\$main
> result.dist\$main.A0
> result.dist\$main.A1
Listing 9: Output matrices of rd2d.dist()
@@
Listing 10: Functionality of rdbw2d.dist()
@@
Listing 11: Functionality of rdbw2d.dist() with kink = "on"

### v2

@@ 2.1.1 Point Estimation, MSE Expansions, and Bandwidth Choices
  • Given a (preliminary) bandwidth, the package rd2d implements several variance estimators for each group $t\in\{0,1\}$: $V_{t,\mathbf{x}}$ is replaced by $\widehat{V}_{t,\mathbf{x}}=\mathbf{e}_{1}^{\top}\widehat{\boldsymbol{\Gamma}}_{t,\mathbf{x}}^{-1}\widehat{\boldsymbol{\Sigma}}_{t,\mathbf{x},\mathbf{x}}\widehat{\boldsymbol{\Gamma}}_{t,\mathbf{x}}^{-1}\mathbf{e}_{1}$, where $\widehat{\boldsymbol{\Sigma}}_{t,\mathbf{x},\mathbf{x}}$ is either a heteroskedasticity-consistent (HC) or cluster-consistent (CR) variance estimator based on replacing the unknown residuals $\varepsilon_{i,t,\mathbf{x}}$ with the plug-in residuals estimates
    $\widehat{\varepsilon}_{i,t,\mathbf{x}}=Y_{i}-\mathbf{R}_{p}(\mathbf{X}_{i}-\mathbf{x})^{\top}\widehat{\boldsymbol{\beta}}_{t}(\mathbf{x}).$
    More precisely, the package allows for the following options:
@@
      Construct an MSE-optimal bandwidth choice for estimating the linear combination given by $\sum_{|\mathbf{k}|=p+1}\frac{\mu_{t}^{(\mathbf{k})}(\mathbf{x})}{\mathbf{k}!}\mathbf{e}_{1}^{\top}\widehat{\boldsymbol{\Gamma}}_{t,\mathbf{x}}^{-1}\widehat{\boldsymbol{\vartheta}}_{t,\mathbf{x}}(\mathbf{k})$, using a $q$th order local polynomial estimator and the bandwidth $\widehat{c}$ for the coefficients $\mathbf{e}_{1}^{\top}\widehat{\boldsymbol{\Gamma}}_{t,\mathbf{x}}^{-1}\widehat{\boldsymbol{\vartheta}}_{t,\mathbf{x}}(\mathbf{k})$. To implement this bandwidth choice, $\widehat{c}$ is used for variance estimation, and a preliminary nearest-neighbor-based polynomial regression approximation is used for bias estimation. The resulting data-driven bandwidth choice is:
@@ 2.1.2 Statistical Inference
where $\widehat{\tau}(\mathbf{x})$ and $\widehat{V}_{\mathbf{x}}$ are constructed using the polynomial order $p$ and bandwidth $h$, after choosing the appropriate HC and CR variance estimator, as explained above in the context of bandwidth selection.
In practice, the bandwidth is chosen to be (I)MSE-optimal, and thus the sampling distribution of the statistic satisfies the following pointwise distributional approximation:
@@
where $\Phi_{\alpha}$ is the $\alpha$th quantile of the standard Gaussian distribution. However, for “large” bandwidths such as the (I)MSE-optimal choice, $\operatorname{CI}(\mathbf{x})$ will be invalid due to the bias $\text{Bias}(\mathbf{x})$, thereby delivering empirical coverage well below its nominal target. A solution to this problem is to employ ad-hoc undersmoothing, that is, to implement $\operatorname{CI}(\mathbf{x})$ with a “smaller” bandwidth $h$ relative to the (I)MSE-optimal choice. Calonico et al. (2018, 2022) showed that undersmoothing is sub-optimal (possibly invalid) under standard assumptions, while the robust bias-correction (RBC) methodology introduced by Calonico et al. (2014) enjoys validity and better (in some cases optimal) higher-order distributional properties. The core idea behind the RBC approach can be summarized as follows: (i) employ the (I)MSE-optimal bandwidth for constructing the point estimator $\widehat{\tau}(\mathbf{x})$, (ii) de-bias (bias correct) the numerator of the statistic $\operatorname{T}(\mathbf{x})$, and (iii) adjust the variance estimate to account for the variability introduced by the debiasing of $\widehat{\tau}(\mathbf{x})$.
@@
which is constructed with an (I)MSE-optimal bandwidth choice for the $p$th order local polynomial point estimator $\widehat{\tau}(\mathbf{x})$. (This amounts to a form of robust bias correction because $\widehat{\tau}_{q}(\mathbf{x})=\widehat{\tau}(\mathbf{x})-h^{p+1}\widehat{B}_{\mathbf{x}}$, where $\widehat{B}(\mathbf{x})$ is an “estimate” of $B_{\mathbf{x}}$.) The package rd2d employs $p=1$ and $q=2$ as defaults, which are standard choices for implementation.
@@
where $\mathsf{q}_{\alpha}$ denotes a suitably chosen quantile to control false rejections uniformly over $\mathcal{B}$. In practice, the continuous assignment boundary is discretized to consider the $J$ cutoff in $\mathbf{b}=(\mathbf{b}_{1},\cdots,\mathbf{b}_{J})^{\top}$ jointly. Then, a feasible quantile choice is
$\mathsf{q}_{\alpha}=\inf\left\{u\geq 0:\ \mathbb{P}\Big{[}\max_{1\leq j\leq J}\big{|}\mathbf{e}_{j}^{\top}\widehat{\mathbf{C}}_{\mathbf{b},q}^{1/2}\mathbf{Z}\big{|}\geq u\Big{|}\text{Data}\Big{]}\leq\alpha\right\},$
where the $J$-dimensional standard Gaussian vector $\mathbf{Z}\thicksim\mathsf{N}(\mathbf{0}_{J},\mathbf{I}_{J})$ is independent of the data, and the $(J\times J)$ covariance matrix $\widehat{\mathbf{C}}_{\mathbf{b},q}=\big{[}\widehat{V}_{\mathbf{b}_{j},\mathbf{b}_{k}}/(\widehat{V}_{\mathbf{b}_{j},\mathbf{b}_{j}}\widehat{V}_{\mathbf{b}_{k},\mathbf{b}_{k}})^{1/2}:1\leq j,k\leq J\big{]}$ where $\widehat{V}_{\mathbf{b}_{j},\mathbf{b}_{k}}=\widehat{V}_{1,\mathbf{b}_{j},\mathbf{b}_{k}}+\widehat{V}_{0,\mathbf{b}_{j},\mathbf{b}_{k}}$ with
@@
$\widehat{\boldsymbol{\Sigma}}_{t,\mathbf{b}_{j},\mathbf{b}_{k}}=\frac{h^{2}}{n}\sum_{i=1}^{n}\mathbf{r}_{q}\Big{(}\frac{\mathbf{X}_{i}-\mathbf{b}_{j}}{h}\Big{)}\mathbf{r}_{q}\Big{(}\frac{\mathbf{X}_{i}-\mathbf{b}_{k}}{h}\Big{)}^{\top}K_{h}(\mathbf{X}_{i}-\mathbf{b}_{j})K_{h}(\mathbf{X}_{i}-\mathbf{b}_{k})\widehat{\varepsilon}_{i,\mathbf{b}_{j}}\widehat{\varepsilon}_{i,\mathbf{b}_{k}}\mathds{1}(\mathbf{X}_{i}\in\mathcal{A}_{t})$
with $\widehat{\varepsilon}_{i,t,\mathbf{x}}=Y_{i}-\mathbf{R}_{q}(\mathbf{X}_{i}-\mathbf{x})^{\top}\widehat{\boldsymbol{\beta}}_{t}(\mathbf{x})$, for $t\in\{0,1\}$. This formulation is robust to unknown conditional heteroskedasticity, while the clustered-robust analogue formula is omitted to save space; see Zeileis (2004) and Zeileis et al. (2020).
@@ 2.1.3 Regularization Strategies
2.2 Aggregated Average Treatment Effects Along the Boundary
For a weight function $w:\mathcal{B}\rightarrow[0,\infty)$, the aggregated average treatment effect (AATE) along the boundary $\mathcal{B}$ is
[…]
A choice of weight function recovers a scalar causal effect, aggregating potential heterogeneous treatment effects along the boundary. Estimation and inference methods for this class of causal parameters can be deduced from our results. For example, consider the generic plug-in “estimator” of $\tau_{\text{AATE},\mathcal{B}}$ given by
$\widehat{\tau}_{\text{AATE},\mathbf{b}}=\frac{\sum_{j=1}^{J}\widehat{\tau}(\mathbf{b}_{j})w(\mathbf{b}_{j})}{\sum_{j=1}^{J}w(\mathbf{b}_{j})},$
where the cutoff points $\mathbf{b}=(\mathbf{b}_{1},\dots,\mathbf{b}_{J})^{\top}$ along the assignment boundary are assumed to be evenly-spaced over $\mathcal{B}$, and $\widehat{\tau}(\mathbf{b}_{j})$ is constructed using $p$th order local polynomial fits. It follows that the IMSE-optimal bandwidth choice is more natural. Feasible RBC confidence intervals are
$\operatorname{CI}_{q}(\text{AATE})=\bigg{[}\widehat{\tau}_{\text{AATE},\mathbf{b},q}\pm\Phi_{1-\alpha/2}\sqrt{\frac{\mathbf{w}^{\top}\widehat{\mathbf{V}}_{\mathbf{b},q}\mathbf{w}}{nh^{2}}}\,\bigg{]},$
where $\widehat{\tau}_{\text{AATE},\mathbf{b},q}$ is the point estimator $\widehat{\tau}_{\text{AATE},\mathbf{b}}$ constructed using a $q$th order polynomial basis, $\widehat{\mathbf{V}}_{\mathbf{b},q}=\big{[}\widehat{V}_{\mathbf{b}_{j},\mathbf{b}_{k}}:1\leq j,k\leq J\big{]}$ denotes its associated variance estimator, as explained previously, and $\mathbf{w}=(w(\mathbf{b}_{1}),\dots,w(\mathbf{b}_{J}))^{\top}/(\sum_{j=1}^{J}w(\mathbf{b}_{j}))$ is the user-chosen vector of weights. In rd2d, AATE estimation and inference are implemented via the optional argument AATE = w in the summary() method for rd2d objects.
@@ 3.1 Function rd2d()
> result.rd2d\$results
> result.rd2d\$results.A0
> result.rd2d\$results.A1
@@
When the optional argument AATE $=\bigl{(}w_{j}\bigr{)}_{j=1}^{J}$, representing the weights for the evaluation points $\{\boldsymbol{b}_{j}\}_{j=1}^{J}$, is supplied to summary(), the output gains an extra row labeled “AATE.” This row reports the aggregated average treatment effect estimator
$\widehat{\tau}_{\text{AATE}}=\frac{\sum_{j=1}^{J}w_{j}\,\widehat{\tau}\!\bigl{(}\boldsymbol{b}_{j}\bigr{)}}{\sum_{j=1}^{J}w_{j}},$
along with its inference statistics.
> summary(result.rd2d, subset = c(1,5,10,15,21,25,30,35,40), AATE = rep(1,neval))
rd2d
Number of Obs. 20000
BW type. mserd-dpi-std
Kernel triangular-prod
VCE method hc1
Masspoints check
Number of Obs. 6191 13809
Estimand (deriv) 0 0
Order est. (p) 1 1
Order rbc. (q) 2 2
Unique Obs. 6191 13809
ID b1 b2 Est. z P>|z| 95% CI
1 0.000 50.000 0.8143 9.8761 0.0000 [0.6191, 0.9257]
5 0.000 40.000 0.7298 11.8383 0.0000 [0.6125, 0.8556]
10 0.000 27.500 0.6752 11.3927 0.0000 [0.5487, 0.7767]
15 0.000 15.000 0.6456 13.4458 0.0000 [0.5448, 0.7307]
21 0.000 0.000 0.6890 8.8464 0.0000 [0.5294, 0.8307]
25 10.000 0.000 0.6647 14.1815 0.0000 [0.5492, 0.7254]
30 22.500 0.000 0.6116 11.5752 0.0000 [0.5049, 0.7107]
35 35.000 0.000 0.5722 8.6064 0.0000 [0.4110, 0.6534]
40 47.500 0.000 0.5120 6.2842 0.0000 [0.3354, 0.6393]
AATE 0.6534 28.9598 0.0000 [0.5987, 0.6856]
Listing 4: Functionality of rd2d() for AATE
@@
1 0.000 50.000 16.687 16.717 16.687 16.717 689 1295
5 0.000 40.000 14.782 14.808 14.782 14.808 1019 1810
10 0.000 27.500 12.736 12.759 12.736 12.759 1252 2108
15 0.000 15.000 12.045 12.067 12.045 12.067 1295 2167
21 0.000 0.000 14.740 14.766 14.740 14.766 2169 1552
25 10.000 0.000 12.455 12.477 12.455 12.477 1467 1980
30 22.500 0.000 11.476 11.496 11.476 11.496 1107 1843
35 35.000 0.000 11.595 11.615 11.595 11.615 824 1418
40 47.500 0.000 14.961 14.987 14.961 14.987 638 1242
Listing 5: Functionality of rd2d() with output = "bw"
@@ 3.2 Function rdbw2d()
The function rdbw2d() is used for MSE (or IMSE) optimal bandwidth implementation, and is used internally in rd2d() when user does not specify bandwidth choices manually. The function takes the same input data as rd2d(), that is, an outcome vector y, a bivariate location score matrix X, a treatment indicator vector t, and a grid of evaluation points b along the boundary $\mathcal{B}$. In addition, the option bwselect encodes four options of bandwidth type: (i) "mserd" finds the MSE-optimal bandwidth for estimating (derivatives of) treatment effect, (ii) "imserd" finds the integreated MSE optimal bandwidth for estimating (derivatives of) treatment effect, (iii) "msetwo" finds the MSE optimal bandwidth for estimation (the derivatives of) conditional means of two potential outcome variables, (iv) "imsetwo" finds the integrated MSE optimal bandwidth for estimation (the derivatives of) conditional means of two potential outcome variables. The default is bwselect = "mserd". An additional Boolean argument stdvar indicates whether the covariates are first standardized to unit standard deviation in each coordinate, in which case the optimal bandwidth is estimated and then converted back to the original scale. The default is stdvar = TRUE.
@@
1 0.000 2.855 20.552 20.588 20.552 20.588
5 0.000 2.284 18.668 18.701 18.668 18.701
10 0.000 1.570 16.476 16.505 16.476 16.505
15 0.000 0.856 14.098 14.123 14.098 14.123
21 0.000 0.000 19.115 19.149 19.115 19.149
25 0.572 0.000 15.737 15.765 15.737 15.765
30 1.287 0.000 14.000 14.025 14.000 14.025
35 2.002 0.000 13.562 13.586 13.562 13.586
40 2.717 0.000 18.798 18.832 18.798 18.832
Listing 6: Functionality of rdbw2d()
@@
30 22.500 0.000 0.6109 7.3862 0.0000 [0.4337, 0.7470]
35 35.000 0.000 0.5768 6.4730 0.0000 [0.3668, 0.6855]
40 47.500 0.000 0.4654 6.7582 0.0000 [0.3602, 0.6544]
Listing 7: Functionality of rd2d.dist()
@@
Listing 8: Functionality of rd2d.dist() with CBuniform = TRUE
When the optional argument AATE $=\bigl{(}w_{j}\bigr{)}_{j=1}^{J}$, representing the weights for the evaluation points $\{\boldsymbol{b}_{j}\}_{j=1}^{J}$, is supplied to summary(), the output gains an extra row labeled “AATE” for estimation and inference of aggregated average treatment effect.
> summary(result.dist, subset = c(1,5,10,15,21,25,30,35,40), AATE = rep(1,neval))
Listing 9: Functionality of rd2d() for AATE
@@
Listing 10: Functionality of rd2d.dist() with output = "bw"
@@
> result.dist\$results
> result.dist\$results.A0
> result.dist\$results.A1
Listing 11: Output matrices of rd2d.dist()
@@
Listing 12: Functionality of rdbw2d.dist()
@@
Listing 13: Functionality of rdbw2d.dist() with kink = "on"